### v1

# Approximation properties of torsion classes

Abstract.
We clarify some results of Bagaria and Magidor [MR3152715] about the relationship between large cardinals and torsion classes of abelian groups, and prove that (1) the Maximum Deconstructibility principle introduced in [Cox_MaxDecon] requires large cardinals; it sits, implication-wise, between Vopěnka’s Principle and the existence of an $\omega_{1}$-strongly compact cardinal. (2) While deconstructibility of a class of modules always implies the precovering property by [MR2822215], the concepts are (consistently) non-equivalent, even for classes of abelian groups closed under extensions, homomorphic images, and colimits.
The first author was supported by the National Science Foundation (DMS-2154141). The second author acknowledges support from the Department of Mathematics at Harvard University as well as from the Harvard Center of Mathematical Sciences and Applications (CMSA). The research of the third author was supported by GAČR 23-05148S
1. Introduction
Classical homological algebra involves approximating an arbitrary module with modules from the well-behaved class $\mathcal{P}_{0}$ of projective modules. Relative homological algebra ([MR178036]) attempts to replace $\mathcal{P}_{0}$ by some other well-behaved class $\mathcal{C}$, leading to new invariants that can be used to create new module and ring theoretic invariants. A crucial requirement for this to work nicely is that $\mathcal{C}$ be a precovering class in the sense of Enochs and Auslander (see Section 2 for definitions).
It can be difficult to show that a class is precovering, but Saorín and Šťovíček proved that
[…]
where deconstructibility is a concept that arose from the solution of the Flat Cover Conjecture ([MR1798574], [MR1832549]) and is closely related to Quillen’s Small Object Argument from homotopy theory. A natural question arose:
Question 1.
Suppose $\mathcal{C}$ is a class of modules that is closed under transfinite extensions. Does the converse of (italic-$\star$1) hold?
After clarifying some results of Bagaria and Magidor from [MR3152715], we use their results to provide a negative answer to Question 1, at least in the absence of large cardinals:
Theorem.
If there are no $\omega_{1}$-strongly compact cardinals, then there is a precovering class of abelian groups that is closed under transfinite extensions, homomorphic images, and colimits, but is not deconstructible.
In fact, the torsion class
${}^{\perp_{0}}\mathbb{Z}:=\{A\in\textbf{Ab}\ :\ \text{Hom}(A,\mathbb{Z})=0\}$
is always a covering class that is closed under transfinite extensions, homomorphic images, and colimits. We prove:
Theorem 2.
${}^{\perp_{0}}\mathbb{Z}$ is deconstructible if and only if there is an $\omega_{1}$-strongly compact cardinal.
Theorem 2 strengthens a result of Bagaria and Magidor, who got the same equivalence with “deconstructible” weakened to “bounded” (we show in Lemma 11 that these concepts are equivalent, for any class of modules closed under transfinite extensions, homomorphic images, and colimits).
In [Cox_MaxDecon], the first author proved that any deconstructible class is:
  1. (a)
    closed under transfinite extensions (by definition), and
  2. (b)
    “eventually almost everywhere closed under quotients”. This is an extremely weak version of saying that if $A\subset B$ are both in the class, then so is $B/A$. In particular, it holds if the class is closed under homomorphic images.
He introduced the Maximum Deconstructiblity principle, which asserts that any class satisfying both (a) and (b) is deconstructible, and proved that
Vopěnka’s Principle (VP) implies Maximum Deconstructibility.
Maximum Deconstructibility appeared very powerful, since [Cox_MaxDecon] showed that it implied deconstructibility of many classes in Gorenstein Homological Algebra that (so far) are not known to be deconstructible in ZFC alone. But it was unclear whether Maximum Deconstructibility had any large cardinal strength at all. We again use the (clarified) results of Bagaria and Magidor to show that it does:
Theorem.
Maximum Deconstructibility implies the existence of an $\omega_{1}$-strongly compact cardinal.
So Maximum Deconstructibility lies, implication-wise, between VP and the existence of an $\omega_{1}$-strongly compact cardinal. The first author still conjectures that it is equivalent to VP.
Section 2 has preliminaries. Section 3 proves the approximation properties of torsion classes that are provable in ZFC alone. Section 4 clarifies some results of Bagaria-Magidor [MR3152715]. Section 5 proves the main theorems mentioned above, and Section 6 includes some questions.
2. Preliminaries
Our notation and conventions follow Kanamori [MR1994835] (for set theory and large cardinals) and Göbel-Trlifaj [MR2985554] (for module theory). By ring we will mean a unital and not necessarily commutative ring. If $R$ is a ring, the class of left $R$-modules will be denoted by $R$-Mod. We will say that $M$ is a module rather than a left $R$-module whenever $R$ is clear from the context. For a regular cardinal $\kappa$ and a module $M$, the collection of all submodules $N$ of $M$ that are ${<}\kappa$-generated will be denoted by $[M]^{<\kappa}$; if $|R|<\kappa$ and $\kappa$ is regular and uncountable, then “${<}\kappa$-generated” is equivalent to “of cardinality less than $\kappa$”. A cardinal $\kappa$ is strongly compact if every $\kappa$-complete filter (on any set whatsoever) can be extended to a $\kappa$-complete ultrafilter. Bagaria and Magidor considered a weakening of strong compactness:
Definition 3 (Bagaria and Magidor, [MR3152715]).
An uncountable cardinal $\kappa$ is called $\omega_{1}$-strongly compact if every $\kappa$-complete filter extends to an $\omega_{1}$-complete ultrafilter.
Note that if $\kappa$ is $\omega_{1}$-strong compact then so is any cardinal $\lambda\geq\kappa$. Unlike strongly compact cardinals, an $\omega_{1}$-strong compact cardinal may fail to be regular [MR3152715, §6].
3. Approximation properties and torsion classes
Given a class $\mathcal{K}$ of $R$-modules, a $\mathcal{K}$-filtration is a $\subseteq$-increasing and $\subseteq$-continuous sequence $\langle M_{\xi}\mid\xi<\eta\rangle$ of modules such that $M_{0}=0$ and for all $\xi<\eta$ such that $\xi+1<\eta$, $M_{\xi}$ is a submodule of $M_{\xi+1}$ and $M_{\xi+1}/M_{\xi}$ is isomorphic to a member of $\mathcal{K}$. A module $M$ is $\mathcal{K}$-filtered whenever there is a $\mathcal{K}$-filtration $\langle M_{\xi}\mid\xi<\eta\rangle$ whose union is $M$. We shall denote the class of all $\mathcal{K}$-filtered modules by $\mathrm{Filt}(\mathcal{K})$. $\mathcal{K}$ is closed under transfinite extensions whenever $\mathrm{Filt}(\mathcal{K})\subseteq\mathcal{K}.$ Finally, $\mathcal{K}^{<\kappa}$ denotes the class of all ${<}\kappa$-presented members of $\mathcal{K}$.
Definition 4.
Let $\mathcal{K}$ be a class of modules and $\lambda\leq\kappa$ be regular cardinals.
  1. (1)
    $\mathcal{K}$ is $\kappa$-deconstructible whenever $\mathcal{K}=\text{Filt}\left({\mathcal{K}}^{<\kappa}\right)$; equivalently, $\mathcal{K}$ is closed under transfinite extensions, and every $M\in\mathcal{K}$ admits a ${\mathcal{K}}^{<\kappa}$-filtration. [^1] This is the definition of deconstructibility in most newer references, such as [sar_trl_DeconAEC], since it is the version that (by [MR2822215]) implies the precovering property. Some other sources (e.g. Cox [Cox_MaxDecon], Göbel-Trlifaj [MR2985554]) only require that $\mathcal{K}\subseteq\text{Filt}\left({\mathcal{K}}^{<\kappa}\right)$ (i.e., with $\subseteq$ rather than equality).
  2. (2)
    $\mathcal{K}$ is $(\kappa,\lambda)$-cofinal if $\mathcal{K}\cap[M]^{<\kappa}$ is $\subseteq$-cofinal in $[M]^{<\lambda}$ for every module $M\in\mathcal{K}$ (i.e., whenever every $N\in[M]^{<\lambda}$ is contained in some $L\in\mathcal{K}\cap[M]^{<\kappa}$).
  3. (3)
    $\mathcal{K}$ is bounded by $\kappa$ whenever $M=\sum\left(\mathcal{K}\cap[M]^{<\kappa}\right)$ for all $M\in\mathcal{K}$ (i.e., if every $x\in M$ is contained in some ${<}\kappa$-presented submodule of $M$ that lies in $\mathcal{K}$). [^2] This property was introduced by Gardner in [MR0647168] and later investigated by Dugas in [MR0799895].
  4. (4)
    $\mathcal{K}$ is $\kappa$-decomposable if every module in $\mathcal{K}$ is a direct sum of ${<}\kappa$-presented modules from $\mathcal{K}$.
$\mathcal{K}$ is said to be deconstructible provided it is $\kappa$-deconstructible for a regular cardinal $\kappa$. The same convention is applied to the rest of above-mentioned properties.
Clearly any $(\kappa,\lambda)$-cofinal class is bounded by $\kappa$. In the forthcoming Lemma 11 we will argue that for certain classes $\mathcal{K}$, the concepts are equivalent.
The following definitions are due to Enochs (generalizing earlier work of Auslander):
Definition 5.
Let $\mathcal{K}$ be a class of modules. We say that $\mathcal{K}$ is:
  1. (1)
    Precovering: If every module $M$ posseses a $\mathcal{K}$-precover; namely, a morphism $f\colon C\rightarrow M$ with $C\in\mathcal{K}$ such that $\mathrm{Hom}(D,f)$ is surjective for all $D\in\mathcal{K}.$
  2. (2)
    Covering: If every module $M$ posseses a $\mathcal{K}$-cover; namely, a $\mathcal{K}$-precover $f\colon C\rightarrow M$ such that for each $g\in\mathrm{End}(C)$ if $fg=f$ then $g$ is an automorphism of $C$.
Saorín and Šťovíček  [MR2822215] proved that if a class $\mathcal{K}$ (of, say, modules) is deconstructible, then it is precovering. In §5 we argue that there are nicely-behaved covering classes which are yet, at least consistently, not deconstructible. The concrete example we have in mind is the torsion class of the abelian group $\mathbb{Z}$ in a context where the set-theoretic universe does not have any $\omega_{1}$-strongly compact cardinals (see Definition 3).
Dickson [MR0191935] introduced the concept of torsion pairs, which are pairs $(\mathcal{A},\mathcal{B})$ such that
$\mathcal{A}={}^{\perp_{0}}\mathcal{B}:=\{X\ :\ \text{Hom}(X,B)=0\text{ for all}B\in\mathcal{B}\}$
and
$\mathcal{B}=\mathcal{A}^{\perp_{0}}:=\{Y\ :\ \text{Hom}(A,Y)=0\text{ for all }A\in\mathcal{A}\}.$
A class of modules is called a torsion class if it is of the form ${}^{\perp_{0}}\mathcal{B}$ for some class $\mathcal{B}$; notice that in this situation,
$\left({}^{\perp_{0}}\mathcal{B},\big{(}{}^{\perp_{0}}\mathcal{B}\big{)}^{\perp_{0}}\right)$
is easily seen to be a torsion pair (the torsion pair cogenerated by $\mathcal{B}$). So torsion classes are exactly those classes that are the left part of some torsion pair. Whenever $\mathcal{X}$ is a singleton $\{X\}$, the convention is to write ${}^{\perp_{0}}X$ rather than {${}^{\perp_{0}}\{X\}$.
By [MR0191935], torsion classes are exactly those classes that are closed under arbitrary direct sums, extensions, and homomorphic images ([MR0389953], Proposition VI.2.1). They have many other desirable features; we list the ones that are relevant for this paper:
Lemma 6.
Torsion classes are:
  1. (1)
    closed under homomorphic images
  2. (2)
    closed under colimits
  3. (3)
    closed under transfinite extensions
  4. (4)
    covering classes.
(1) is immediate from the definition of torsion class, and (2) follows from closure under direct sums and homomorphic images. Closure under transfinite extensions follows easily from closure under extensions and colimits. Finally, given any module $M$ and any torsion class $\mathcal{T}:={}^{\perp_{0}}\mathcal{X}$, the trace of $\mathcal{T}$ in $M$ is defined as
$r(M):=\sum_{T\in\mathcal{T}}\{\text{im}\ \pi\ :\ \pi\in\text{Hom}(T,M)\}.$
Then it is easily seen that $r(M)\in\mathcal{T}$ and that the inclusion $r(M)\to M$ is a $\mathcal{T}$-cover of $M$.
4. Clarification of some results of Bagaria and Magidor
In Bagaria-Magidor [MR3152715, §5] a torsion class ${}^{\perp_{0}}\mathcal{X}$ is called “$\kappa$-generated” whenever every $A\in{}^{\perp_{0}}\mathcal{X}$ is a direct sum of subgroups in ${}^{\perp_{0}}\mathcal{X}$, each of cardinality ${<}\kappa$; this is what is usually called $\kappa$-decomposable (see Definition 4). However, their use of the word “direct” on page 1867 of [MR3152715] appears to be a misprint, since the arguments there never use (or conclude) directness of the relevant sums. All of the arguments in [MR3152715, §5], and the citation they provide for the concept (Dugas [MR0799895]), appear to use the weaker property that every group in the class is merely a sum of subgroups in the class of cardinality ${<}\kappa$. The latter condition is what we called bounded by $\kappa$ in Section 2 (this terminology was used by Gardner [MR0647168] and Dugas [MR0799895]).
The proof of [MR3152715, Theorem 5.1] shows that if $\kappa$ is a $\omega_{1}$-strongly compact cardinal and $X$ is a countable abelian group, then ${}^{\perp_{0}}X$ is bounded by $\kappa$. In fact, they prove the stronger property of ${}^{\perp_{0}}X$ being $(\kappa,\omega_{1})$-cofinal. [^3] Recall that this stands for the following property: For each $G\in{}^{\perp_{0}}\mathbb{Z}$ every countable subgroup $H\leq G$ is included in a member of ${}^{\perp_{0}}\mathbb{Z}\cap[G]^{<\kappa}$ (see Definition 4). The proof however does not show that ${}^{\perp_{0}}X$ is $\kappa$-decomposable, and in fact that is impossible, because of the following theorem (due to the third author):
Theorem 7.
${}^{\perp_{0}}\mathbb{Z}$ is not decomposable.
Proof.
Suppose toward a contradiction that ${}^{\perp_{0}}\mathbb{Z}$ is $\kappa$-decomposable, with $\kappa$ (without loss of generality) a regular uncountable cardinal. Fix any prime $p$. Then all $p$-groups [^4] Abelian groups whose elements all have order a power of $p$. are in ${}^{\perp_{0}}\mathbb{Z}$. So, in particular, the $\kappa$-decomposability assumption of ${}^{\perp_{0}}\mathbb{Z}$ implies
[…]
$P_{\kappa^{+}}=\bigoplus_{i\in I}Q_{i}$
for some collection $(Q_{i})_{i\in I}$ of $<\kappa$-sized subgroups. Since subgroups of $p$-groups are also $p$-groups, each $Q_{i}$ has a $p$-length, which is $<\kappa$ because $|Q_{i}|<\kappa$. And it is easy to check that the $p$-length of a direct sum is at most the supremum of the $p$-lengths of the direct summands, so $\bigoplus_{i\in I}Q_{i}$ has $p$-length at most $\kappa$, contradicting that the $p$-length of $P_{\kappa^{+}}$ is $\kappa^{+}+1$. ∎
The proof of their Theorem 5.1 really showed:
Theorem 8 (clarification of [MR3152715, Theorem 5.1]).
If $\kappa$ is a $\delta$-strongly compact cardinal and $X$ is an abelian group of cardinality less than $\delta$, then ${}^{\perp_{0}}X$ is bounded in $\kappa$.
In the other direction, Theorem 5.3 of [MR3152715] asserted that if ${}^{\perp_{0}}\mathbb{Z}$ is “$\kappa$-generated” (using their definition, with “direct sum” instead of just “sum”), then $\kappa$ is $\omega_{1}$-strongly compact. Fortunately, their argument never used the “direct” part of their definition, since by Theorem 7, that would have been an inconsistent assumption. Their proofs really showed:
Theorem 9 (clarification of [MR3152715, Theorem 5.3]).
If ${}^{\perp_{0}}\mathbb{Z}$ is bounded in $\kappa$, then $\kappa$ is $\omega_{1}$-strongly compact.
Corollary 10 (clarification of [MR3152715, Corollary 5.4]).
${}^{\perp_{0}}\mathbb{Z}$ is bounded in $\kappa$ if and only if $\kappa$ is $\omega_{1}$-strongly compact.
5. Maximum Deconstructibility, and deconstructibility of torsion classes, require large cardinals
Through this section $\kappa$ will denote a regular uncountable cardinal. Our main goal is to clarify, for a given class of modules $\mathcal{K}$, the relationship between $\mathcal{K}$ being $\kappa$-deconstructible and $\mathcal{K}$ being bounded in $\kappa$. As noted in §2, any $(\kappa,\kappa)$-cofinal class $\mathcal{K}$ is bounded in $\kappa$. In fact the former property seems to be strictly stronger than the latter, and they both follow from $\kappa$-deconstructibility. Lemma 11 provides some important scenarios where these concepts to be equivalent. Combining this lemma with the (clarified) results of Bagaria and Magidor from Section 4 allows us to prove the main results mentioned in the introduction, namely:
  1. (1)
    that the Maximum Deconstructibility principle introduced in [Cox_MaxDecon] entails the existence of large cardinals;
  2. (2)
    that in the absence of $\omega_{1}$-strongly compact cardinals, the class ${}^{\perp_{0}}\mathbb{Z}$ is a covering class (cf. Lemma 6 above) that is not deconstructible.
Lemma 11.
Suppose $\kappa$ is a regular uncountable cardinal and $R$ is a ring of size less than $\kappa$, and $\mathcal{K}$ is a class of $R$-modules. Consider the following statements:
  1. (a)
    $\mathcal{K}$ is $\kappa$-deconstructible.
  2. (b)
    $\mathcal{K}$ is $(\kappa,\kappa)$-cofinal and closed under transfinite extensions.
  3. (c)
    $\mathcal{K}$ is bounded by $\kappa$ and closed under transfinite extensions.
Then:
  1. (1)
    (a) $\implies$ (b) $\implies$ (c).
  2. (2)
    If $\mathcal{K}$ is closed under quotients—i.e., if
    $\big{(}A\subset B,\ A\in\mathcal{K},\text{ and }B\in\mathcal{K}\big{)}\implies\ B/A\in\mathcal{K}$
    for all modules $A$ and $B$—then (a) and (b) are equivalent.
  3. (3)
    If $\mathcal{K}$ is closed under homomorphic images and colimits, then all three statements are equivalent.
Proof.
The (a) $\Rightarrow$ (b) direction follows from the Hill Lemma (cf. Göbel-Trlifaj [MR2985554], Theorem 7.10). The (b) $\Rightarrow$ (c) implication is immediate from the definitions.
Now suppose $\mathcal{K}$ is closed under transfinite extensions and quotients. We prove (b) $\Rightarrow$ (a). Without losing any generality we may assume that (the universe of) every module $M\in\mathcal{K}$ is a cardinal. Since $\mathcal{K}$ is closed both under transfinite extensions and quotients, then by [Cox_MaxDecon, Theorem 6.3] in order to prove $\kappa$-deconstructibility of $\mathcal{K}$ it suffices to show that if $M\in\mathcal{K}$, then $\mathcal{K}\cap\wp^{*}_{\kappa}(M)$ is stationary in $\wp^{*}_{\kappa}(M)$, where $\wp^{*}_{\kappa}(M)$ denotes the ${<}\kappa$-sized submodules of $M$ whose intersection with $\kappa$ is transitive.
The set $\mathcal{K}\cap\wp^{*}_{\kappa}(M)$ being stationary in $\wp^{*}_{\kappa}(M)$ is equivalent to the following statement: For each algebra $\mathfrak{A}=(M,\dots)$ on $M$ in a countable signature, there is $X\prec\mathfrak{A}$ in $\mathcal{K}\cap\wp^{*}_{\kappa}(M)$. We prove this next.
Using the assumption that $\mathcal{K}$ is $\subseteq$-cofinal in $[M]^{<\kappa}$, and the downward Löwenheim-Skolem Theorem, we construct an $\subseteq$-increasing $\omega$-chain
$X_{0}\subseteq C_{0}\subseteq X_{1}\subseteq C_{1}\subseteq\dots$
such that for each $n\in\omega$:
  • $X_{n}\prec\mathfrak{A}$, $|X_{n}|<\kappa$, and $X_{n}$ has transitive intersection with $\kappa$.
  • $C_{n}\in\mathcal{K}$ and $|C_{n}|<\kappa$ (but might have non-transitive intersection with $\kappa$).
Let $X:=\bigcup_{n<\omega}X_{n}$. Then $X\prec\mathfrak{A}$, $|X|<\kappa$, and $X\cap\kappa$ is transitive. Also, note that $X=\bigcup_{n<\omega}C_{n}$ too. Since $\mathcal{K}$ is closed under quotients, each $C_{n+1}/C_{n}$ is in $\mathcal{K}$. So since $\mathcal{K}$ is closed under transfinite extensions, and each $C_{n+1}/C_{n}$ is in $\mathcal{K}$, we conclude that $X\in\mathcal{K}$.
Now suppose that $\mathcal{K}$ is closed under homomorphic images and colimits; we show that (c) implies (b) (and hence (a) will hold too, since closure under homomorphic images trivially implies closure under quotients in the sense of 2). Suppose $\mathcal{K}$ is bounded in $\kappa$. Suppose $M\in\mathcal{K}$, and fix any $X\in[M]^{<\kappa}$. By $\kappa$-boundedness of $\mathcal{K}$, for each $x\in X$ there is a $Y_{x}\in\mathcal{K}\cap[M]^{<\kappa}$ with $x\in Y_{x}$. Then $S:=\sum_{x\in X}Y_{x}$ is a $<\kappa$-sized submodule of $M$ and contains $X$. And $S$ is a homomorphic image of the colimit of the (possibly non-directed) diagram of inclusions
$\big{\{}Y_{x}\to Y_{z}\ :\ Y_{x}\subseteq Y_{z}\text{ and }x,z\in X\big{\}}.$
So by closure of $\mathcal{K}$ under homomorphic images and colimits, $S\in\mathcal{K}$. ∎
We now focus on abelian groups. Recall that by Lemma 6 any torsion class ${}^{\perp_{0}}\mathcal{X}$ is closed under homomorphic images, transfinite extensions, and colimits. So by Lemma 11, a class of the form ${}^{\perp_{0}}\mathcal{X}$ is deconstructible if and only if it is bounded in some cardinal. Furthermore, as noted in the introduction, although an $\omega_{1}$-strongly compact cardinal might be singular, all cardinals above it are also $\omega_{1}$-strongly compact. In particular (by considering its successor if necessary), there exists an $\omega_{1}$-strongly compact cardinal if and only if there exists a regular $\omega_{1}$-strongly compact cardinal. Then Lemma 6, Lemma 11, and the (clarified) Corollary 10 of Bagaria and Magidor immediately imply Theorem 2 from the introduction, which asserted that ${}^{\perp_{0}}\mathbb{Z}$ is deconstructible if and only if there is an $\omega_{1}$-strongly compact cardinal. In fact, making use of the (clarified) Theorem 8 of Bagaria and Magidor, we have:
Corollary 12.
The following are equivalent:
  1. (1)
    There exists an $\omega_{1}$-strongly compact cardinal.
  2. (2)
    ${}^{\perp_{0}}\mathbb{Z}$ is deconstructible.
  3. (3)
    For all countable abelian groups $X$, ${}^{\perp_{0}}X$ is deconstructible.
Since ${}^{\perp_{0}}\mathbb{Z}$ is closed under transfinite extensions and homomorphic images, we solve part of [Cox_MaxDecon, Question 8.2] in the affirmative:
Corollary 13.
Maximum Deconstructibility of [Cox_MaxDecon] has large cardinal consistency strength. Specifically, it lies implication-wise, in between the existence of an $\omega_{1}$-strongly compact and Vopěnka’s Principle.
Recall that deconstructible classes of modules are always precovering ([MR2985554, Theorem 7.21]). Thus, another interesting corollary of Corollary 12 and Lemma 6 is the existence of precovering classes (even covering classes) which are not deconstructible:
Corollary 14.
If there are no $\omega_{1}$-strongly compact cardinals, then there is a covering class in Ab that is closed under colimits, transfinite extensions, and homomorphic images, but is not deconstructible (namely, the class ${}^{\perp_{0}}\mathbb{Z}$). Thus, $\mathrm{ZFC}$ cannot prove that all subclasses of Ab satisfying the clauses in Lemma 6 are deconstructible.∎
6. Questions
Question 15.
Suppose that $\kappa\geq\omega_{2}$ is a regular cardinal such that ${}^{\perp_{0}}X$ is deconstructible for all abelian group of size ${<}\kappa$. Must $\kappa$ be strongly compact?
Much of the literature focuses on deconstructibility and precovering properties of “roots of Ext” classes, i.e., classes of the form
:={${}^{\perp}\mathcal{B}:=\big{\{}A\ :\ \forall B\in\mathcal{B}\ \text{Ext}^{1}(A,B)=0\big{\}}$
for some fixed class $\mathcal{B}$ of modules. Our results in Section 5 consistently separates deconstructibility from precovering for ${}^{\perp_{0}}\mathbb{Z}$, but this is very far from being a root of Ext; it does not even contain the ring $\mathbb{Z}$.
This suggests asking whether deconstructibility is equivalent to precovering for classes of the form ${}^{\perp}\mathcal{B}$, but there is a caveat: the first author proved in [Cox_VP_CotPairs] that it is consistent, relative to consistency of Vopěnka’s Principle (VP), that over every hereditary ring (such as $\mathbb{Z}$), every class of the form ${}^{\perp}\mathcal{B}$ is deconstructible. So deconstructibility is trivially equivalent to precovering for roots of Ext (over hereditary rings) in that model. But the following questions are open:
Question 16.
Is it consistent with ZFC that there is a class $\mathcal{B}$ of abelian groups such that ${}^{\perp}\mathcal{B}$ is precovering, but not deconstructible?
Question 17.
Does ZFC prove the existence of a ring $R$ and a class $\mathcal{B}$ of $R$-modules, such that ${}^{\perp}\mathcal{B}$ is precovering, but not deconstructible? By the remarks above, such a ring could not be provably hereditary (unless VP is inconsistent).
Both questions are even open if we replace “classes of the form ${}^{\perp}\mathcal{B}$” with “classes that are closed under transfinite extensions and contain the ring” (such classes would contain all free modules).